A box of contains three oranges and 5 cherry candies. If 4 candies are chosen at random what is the probability that no more than one will be orange?
Answer by stanbon(75887)

A box of contains three oranges and 5 cherry candies. If 4 candies are chosen at random what is the probability that no more than one will be orange?
----
# of ways to succeed: 3C0*5C4 + 3C1*5C3 = 1*5+3*10 = 35
---
# of possible outcomes: 8C4 = (8*7*6*5)/(1*2*3*4) = 70
----
P(0<= x <=1 orange) = 35/70 = 1/2
-------------------------------------
Cheers,
Stan H.